OpenStudy (anonymous):

solve for x: 7x+2y=5

OpenStudy (lgbasallote):

subtract 2y from both sides 7x = 5 - 2y divide both sides by 7 x = (5-2y)7 does that help?

OpenStudy (lgbasallote):

that should be x = (5-2y)/7

OpenStudy (anonymous):

I think so

OpenStudy (lgbasallote):

@ashna it's solve for x

OpenStudy (anonymous):

I am totally confused now

OpenStudy (lgbasallote):

@rannsan just follow what i said above

OpenStudy (anonymous):

7x+2y=5 Subtract the 2y on both sides 7x=-2y+5 Now divided both sides of the equation by 7 x=(-2/7)y+5/7

OpenStudy (anonymous):

x=(-2/7)y+5/7 ?

OpenStudy (anonymous):

I still am not following

OpenStudy (anonymous):

Let us go step by step: Subtract 2y both the sides first.. Can you do this @rannsan

OpenStudy (anonymous):

trying, this is my first algebra class and it has totally confused me

OpenStudy (anonymous):

See I give you similar example: Suppose we have: 5x+4y=10 Now subtract 4y both the sides:  Now to find x we must get rid of 5 attached to it so divide by 5 both the sides:  Getting @rannsan